OpenStudy (anonymous):

A train traveling 50 mph left a station 30 minutes before a second train running at 55 mph. How soon did the second train overtake the first? If 50x represents the distance the slower train travels, then the faster train travels A.55(x - 0.5) B.55(x + 0.5) C.50 + 30x

OpenStudy (anonymous):

A train traveling 50 mph left a station 30 minutes before a second train running at 55 mph. How soon did the second train overtake the first? --------------------------- 1st train DATA: rate = 50 mph ; time = x hrs ; distance = rt = 50x miles ------------------------------ 2nd train DATA: rate = 55 mph ; time = (x-(1/2)) hrs ; distance = rt = 55(x-(1/2)) miles ---------------------------- Equation: distance = distance 50x = 55(x-(1/2) 50x = 55x - (55/2) -5x = -55/2 x = 11/2 x = 5 1/2 hrs (time at which the 2nd train overtakes the 1st train) is A if I deserve a medal than click good answer,Thank

OpenStudy (anonymous):

Thanks Mrs. Lam

OpenStudy (anonymous):

urwc